# SAMM: Sharded Automated Market Makers

### v1

@@ Appendix G Equilibrium Proof Details
$\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}R^{A}_{i^{\ast}},b^{\textit{BA}}_{j})\leq\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}}_{j})\,\,.$ (65)
@@
G.2 Proof of Theorem 7.1.3
  • Theorem 7.1.3 (restated). In $\Gamma_{n}(tf_{\textit{SAMM}})$, considering the following dominant strategy of the BA trader which randomly selects one of the smallest pools to acquire all required tokens:
@@
$U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},\tau^{BA})=\sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})}\frac{1}{n_{\min}(\textbf{R})}\times U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})\$ (69)
The revenue of the BA trader under action $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})$ is (using Equation 6.1.4)
@@
Combining Equation 69 and 70, we find the utility of the BA trader with strategy $\tau^{BA}$
@@
$=$ $\sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})}\frac{1}{n_{\min}(\textbf{R})}\times\left(-p^{B}\times G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{B}_{\min},b^{\textit{BA}})\right)$
@@
$\sum_{i=1}^{n}\left(G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}}_{i})\right)>G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}})\,\,.$
@@
Combining the conditions for Equation 76 and 78, then $\forall\tilde{a}^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, we have
@@
From lemma 7.1.1, the revenue of any action ${a}_{t}^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})$ is no larger than the revenue of the action $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})$. Combining Equation 70, we have
@@
Combining Equations 79 anc 81, we expand the utility of the BA trader under strategy $\pi^{BA}$ in Equation G.2 as
@@
G.3 Proof of Lemma 7.2.1
  • Lemma 7.2.1 (restated). For any action of liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$, if $\rho^{A}(\textbf{R}+a_{lp})\geq\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})$, then $a_{lp}=a_{lp}^{\textit{fill}}$.
Proof.
Consider any $j$ s.t. $l^{A^{\ast}}_{j}>0$, from the definition of $a_{lp}^{\textit{fill}}$ (Definition 7.2.1), $l^{A^{\ast}}_{j}+R^{A}_{j}$ is the minimal among all pools with state $\textbf{R}+a_{lp}$, i.e.,
@@
Combining Expressions 84 8586 and 87, all inequation holds with equality, which indicates that $a_{lp}=a_{lp}^{\textit{fill}}$:
$\forall 1\leq i\leq n,l^{A}_{i}=\hat{l}^{A}_{i}\,\,.\qed$ (88)
G.4 Proof of Lemma 7.2.1
  • Lemma 7.2.1 (restated). For any two pools $i$ and $j$, if $R^{A}_{i}<R^{A}_{j}$, for any output amount $b^{\textit{BA}}$ of token B, the trading fee of $\textit{pool}_{i}$ is strictly smaller than the trading fee of $\textit{pool}_{j}$:
@@
Combining Equation 89 and 90, we conclude that the trading fee of $\textit{pool}_{i}$ is strictly smaller than the trading fee of $\textit{pool}_{j}$:
$tf_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}})<tf_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})\,\,.\qed$ (91)
G.5 Proof of Theorem 7.2.2
  • Theorem 7.2.2 (restated). Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
@@
Proof.
@@
Consider any action of liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$, the pool state after this action is $\textbf{R}+a_{lp}$. From Corollary 7.1.3, any best response of the trader only has a positive probability of taking an action that trades in exactly one of the smallest pools. Therefore, from Equation 18, when the trader use any best response $\tau^{BA}$, the liquidity provider’s revenue with action $a_{lp}$ is
@@
From Equation 6.1.4, the revenue of a liquidity provider due to action $a_{lp}$ and the BA trader action $a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}+a_{lp})$ is acquiring all $b^{\textit{BA}}$ in one of the smallest pool after the liquidity provider’s action, say $\textit{pool}_{i}$, is
@@
From Lemma 7.2.1, the trading fee of the smallest pool is no larger than the trading fee of any other pool:
@@
Combining Equation 96 and 97, we have
@@
Combining Equation 92 and 99, the revenue of the liquidity provider under action $\hat{a}_{lp}$ is not smaller than the revenue of the liquidity provider under any other action when traders use any best response $\tau^{BA}$.
@@
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\left(\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},l^{A},l^{B},a_{lp},\tau^{BA},\tau^{AB})\right)$
@@
G.6 Proof of Theorem 7.2.3
  • Theorem 7.2.3 (restated). In $\Gamma_{n}(\textit{tf}_{\textit{SAMM}})$, in all SPNE, the liquidity provider’s best response is the fillup strategy:
@@
Proof.
@@
(1) Construction of the new action: Consider any action of a liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$. We construct another action $a_{lp}^{\prime}$. First, we want the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ to be larger than the smallest pool in $\textbf{R}+a_{lp}$, which lead to a higher trading fee in a single trade according to Lemma 7.2.1. Second, we want the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ to be unique and also the smallest in R to make the liquidity provider have more shares in the smallest pool than in $\textbf{R}+a_{lp}$.
@@
Consider the fillup action $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})=\left(\left(\hat{l}^{A}_{1},\hat{l}^{B}_{1}\right),\cdots,\left(\hat{l}^{A}_{n},\hat{l}^{B}_{n}\right)\right)$. If $a_{lp}=a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$, we take $a_{lp}^{\prime}=a_{lp}$ and we are done. If $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$, from Lemma 7.2.1, we have $\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp})<\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}))$. Denote this difference by $z=\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}))-\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp})$. We construct the action $a_{lp}^{\prime}=\left(\left(l^{A^{\prime}}_{1},l^{B^{\prime}}_{1}\right),\cdots,\left(l^{A^{\prime}}_{n},l^{B^{\prime}}_{n}\right)\right)$ as follows. Figure 2 shows an example of the construction.
@@
From the Definition 7.2.1, $\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})=l^{A^{\ast}}_{i^{\ast}}+R^{A}_{i^{\ast}}$. Then, the smallest amount of deposited token A in $\textbf{R}+a_{lp}^{\prime}$ is larger than that in $\textbf{R}+a_{lp}$:
@@
(2) The revenues of actions $a_{lp}$ and $a_{lp}^{\prime}$: To compare the revenue due to both actions, we consider the strategies and utility of the subsequent trader. Any best response of the trader $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ uses the smallest pool (Lemma 7.1.3):
@@
Thus, the revenue following action $a_{lp}^{\prime}$ is (using Equation 6.1.4):
@@
Interpreting Equation G.6, the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ is larger than the smallest pool in $\textbf{R}+a_{lp}$:
@@
From lemma 7.2.1, the trading fee of trading in a larger pool is larger,
$\textit{tf}(R^{A}_{i^{\ast}}+l^{A^{\prime}}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+l^{A^{\prime}}_{i^{\ast}}),b^{\textit{BA}})>\textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\,\,.$ (114)
From the definition of $i^{\ast}$ in Equation G.6, we have
@@
Combine Equations 113 and 115, we have
@@
Combining Equations 110111114 and 116, the revenue of the liquidity provider with action $a_{lp}^{\prime}$ is higher than with action $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$:
@@
Combining Equations 108 and 109, the utility of the liquidity provider under action $a_{lp}^{\prime}$ is higher than that under action $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$:
$U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})>U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (118)
@@
$U_{lp}(\textbf{R},\pi_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})=\\ \sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (120)
@@
$=$ $\sum_{a_{lp}^{\prime}\in\mathcal{A}_{lp}(l^{A},l^{B})}\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},a_{lp}^{\prime})\times U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})$
@@
Combining Equations 120  G.6122 and 123, the utility of the liquidity provider under $\pi_{lp}^{\prime}$ is higher than that under $\pi_{lp}$:
$U_{lp}(\textbf{R},\pi_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})>U_{lp}(\textbf{R},\pi_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$

### v2

@@ G-A1 Traders’ optimal action
$\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}R^{A}_{i^{\ast}},b^{\textit{BA}}_{j})\leq\\ \sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}}_{j})\,\,.$ (63)
@@
G-A2 Using smallest pools is a dominant strategy
Lemma G-A1 indicates that when multiple AMM pools have the same smallest amount of deposited tokens, acquiring all tokens in any one of them has the highest utility. Since the utility of a trader’s strategy is the linear combination of the utility of actions, it is a dominant strategy for the trader to randomly select one of the smallest pools to acquire all required tokens: {cor}[] In $\Gamma_{n}(tf_{\textit{SAMM}})$, a dominant strategy for a BA trader is to randomly select one of the smallest pools to acquire all required tokens:
If $n_{\min}(\textbf{R})=n$, then all pools have the same amount of deposited tokens, and the trader randomly selects one of the $n$ pools. {cor}[] In $\Gamma_{n}(tf_{\textit{SAMM}})$, the system state is R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$. If $\forall i,j,R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, then the perfect parallelism strategy
is a dominant strategy for the BA trader.
G-A3 All dominant strategies use smallest pools
We have shown that only trading in one of the smallest pools is the dominant strategy for the trader. However, to later determine the best response of liquidity providers, we need to know whether there are other dominant strategies. We show that if the trader has a positive probability of taking the action of splitting a transaction or trading in a pool with not the smallest amount of deposited tokens, then she has strictly lower utility than randomly selecting one of the smallest pools to trade, as we intended:
  • Theorem VI-C1 (restated). In $\Gamma_{n}(tf_{\textit{SAMM}})$, considering the following dominant strategy of the BA trader which randomly selects one of the smallest pools to acquire all required tokens:
@@ G-A3 All dominant strategies use smallest pools
$U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},\tau^{BA})=\\ \sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})}\frac{1}{n_{\min}(\textbf{R})}\times U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})\$ (67)
The revenue of the BA trader under action $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})$ is (using Equation VI-A4)
@@
Combining Equation 67 and 68, we find the utility of the BA trader with strategy $\tau^{BA}$
@@
$=$ $\sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})}\frac{\left(-p^{B}\times G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{B}_{\min},b^{\textit{BA}})\right)}{n_{\min}(\textbf{R})}$
@@
$\sum_{i=1}^{n}\left(G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}}_{i})\right)>\\ G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}})\,\,.$
@@
Combining the conditions for Equation 74 and 76, then $\forall\tilde{a}^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, we have
@@
From lemma G-A1, the revenue of any action ${a}_{t}^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})$ is no larger than the revenue of the action $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})$. Combining Equation 68, we have
@@
Combining Equations 77 anc 79, we expand the utility of the BA trader under strategy $\pi^{BA}$ in Equation G-A3 as
@@
From the above theorem, all the best responses of the trader should only have a positive probability of taking an action that trades in exactly one of the smallest pools: {cor}[] Considering any best response strategy $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ of the BA trader, the strategy should only have a positive probability of taking an action that trades in one of the smallest pools:
G-B Liquidity Provider Strategy and SPNE
Our analysis of the trader strategy shows that if the shards are balanced, traders behave as intended. Next, we consider the strategies of liquidity providers in equilibrium. They should maintain the shard balance. Moreover, their incentives should rebalance the system state even in the face of attacks that break this balance.
G-B1 Scaffolding
We want a liquidity provider to fill up smaller pools to keep pools balanced. We call such an action the fillup action, where if the liquidity provider inputs tokens in a pool, then the pool is the smallest pool after this action. We denote the fillup action by $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$: {defi}[] The fillup action of a liquidity provider $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})=\left(\left(\hat{l}^{A}_{1},\hat{l}^{B}_{1}\right),\cdots,\left(\hat{l}^{A}_{n},\hat{l}^{B}_{n}\right)\right)$ is the action satisfying that if the liquidity provider inputs tokens in a pool, then the pool is one of the smallest pools after this action:
$\forall 1\leq i\leq n:\hat{l}^{A}_{i}\geq 0,$
$\sum_{i=1}^{n}\hat{l}^{A}_{i}=l^{A}\,\,;$
We also define the strategy that only takes the fillup action as the fillup strategy: {defi}[] The fillup strategy of a liquidity provider $\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B})$ is the strategy that only takes the fillup action:
$\rho^{A}(\textbf{R})=\min_{1\leq i\leq n}R^{A}_{i}\,.$
{lem}
[] For any action of liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$, if $\rho^{A}(\textbf{R}+a_{lp})\geq\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})$, then $a_{lp}=a_{lp}^{\textit{fill}}$.
Proof:
Proof:
Consider any $j$ s.t. $l^{A^{\ast}}_{j}>0$, from the definition of $a_{lp}^{\textit{fill}}$ (Definition G-B1), $l^{A^{\ast}}_{j}+R^{A}_{j}$ is the minimal among all pools with state $\textbf{R}+a_{lp}$, i.e.,
@@ G-B1 Scaffolding
Combining Expressions 83 8485 and 86, all inequation holds with equality, which indicates that $a_{lp}=a_{lp}^{\textit{fill}}$:
$\forall 1\leq i\leq n,l^{A}_{i}=\hat{l}^{A}_{i}\,\,.$ (87)
We have shown that traders are incentivized to trade in smaller pools in Section G-A, which incentivizes liquidity providers to input their tokens in smaller pools. Additionally, if the liquidity provider makes small pools larger, she would get more trading fees, which further incentivizes them to add liquidity to small pools:
{lem}
[] For any two pools $i$ and $j$, if $R^{A}_{i}<R^{A}_{j}$, for any output amount $b^{\textit{BA}}$ of token B, the trading fee of $\textit{pool}_{i}$ is strictly smaller than the trading fee of $\textit{pool}_{j}$:
@@
Combining Equation 88 and 89, we conclude that the trading fee of $\textit{pool}_{i}$ is strictly smaller than the trading fee of $\textit{pool}_{j}$:
$tf_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}})<tf_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})\,\,.$ (90)
G-B2 Perfect parallelism under balanced pools
  • Theorem VI-C2 (restated). Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
@@ G-B2 Perfect parallelism under balanced pools
Proof:
Traders prefer trading in smaller pools to reduce their costs. At the same time, fees are higher in larger pools. So the liquidity provider should increase her share in the smallest pools. This dual objective is optimally achieved by uniformly distributing tokens across all pools. ∎
Proof:
@@
Consider any action of liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$, the pool state after this action is $\textbf{R}+a_{lp}$. From Corollary G-A3, any best response of the trader only has a positive probability of taking an action that trades in exactly one of the smallest pools. Therefore, from Equation 18, when the trader use any best response $\tau^{BA}$, the liquidity provider’s revenue with action $a_{lp}$ is
@@
From Equation VI-A4, the revenue of a liquidity provider due to action $a_{lp}$ and the BA trader action $a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}+a_{lp})$ is acquiring all $b^{\textit{BA}}$ in one of the smallest pool after the liquidity provider’s action, say $\textit{pool}_{i}$, is
@@
From Lemma G-B1, the trading fee of the smallest pool is no larger than the trading fee of any other pool:
@@
Combining Equation 95 and 96, we have
@@
Combining Equation 91 and 98, the revenue of the liquidity provider under action $\hat{a}_{lp}$ is not smaller than the revenue of the liquidity provider under any other action when traders use any best response $\tau^{BA}$.
@@
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\left(\begin{subarray}{c}\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\\ \times U_{lp}(\textbf{R},l^{A},l^{B},a_{lp},\tau^{BA},\tau^{AB})\end{subarray}\right)$
@@
{cor}
[] Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
constitute an SPNE.
G-B3 Convergence to Balanced Pools
We now show that even if the system reaches an unbalanced state, maybe due to an attacker, it converges to a balanced state since the liquidity provider uses the fillup strategy. We can conclude that the fillup strategy is the only best response in all SPNE:
  • Theorem VI-C2 (restated). In $\Gamma_{n}(\textit{tf}_{\textit{SAMM}})$, in all SPNE, the liquidity provider’s best response is the fillup strategy:
@@ G-B3 Convergence to Balanced Pools
Proof:
Given any action $a_{lp}$ that is not the fillup action, we can construct a new action $a_{lp}^{\prime}$ that is strictly better than $a_{lp}$. By ensuring the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ is larger than that in $\textbf{R}+a_{lp}$, we increase trading fees garnered from each transaction. Moreover, this smallest pool is also the smallest before the action, maximizing the liquidity provider’s share. Consequently, the liquidity provider earns higher revenue under $a_{lp}^{\prime}$ than under $a_{lp}$. Thus, any strategy incorporating an action other than the fillup action is not optimal. Figure 3 illustrates an example o  $a_{lp}^{\prime}$ construction. ∎
Proof:
@@
(1) Construction of the new action: Consider any action of a liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$. We construct another action $a_{lp}^{\prime}$. First, we want the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ to be larger than the smallest pool in $\textbf{R}+a_{lp}$, which lead to a higher trading fee in a single trade according to Lemma G-B1. Second, we want the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ to be unique and also the smallest in R to make the liquidity provider have more shares in the smallest pool than in $\textbf{R}+a_{lp}$.
@@
Consider the fillup action $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})=\left(\left(\hat{l}^{A}_{1},\hat{l}^{B}_{1}\right),\cdots,\left(\hat{l}^{A}_{n},\hat{l}^{B}_{n}\right)\right)$. If $a_{lp}=a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$, we take $a_{lp}^{\prime}=a_{lp}$ and we are done. If $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$, from Lemma G-B1, we have $\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp})<\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}))$. Denote this difference by $z=\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}))-\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp})$. We construct the action $a_{lp}^{\prime}=\left(\left(l^{A^{\prime}}_{1},l^{B^{\prime}}_{1}\right),\cdots,\left(l^{A^{\prime}}_{n},l^{B^{\prime}}_{n}\right)\right)$ as follows. Figure 3 shows an example of the construction.
@@
From the Definition G-B1, $\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})=l^{A^{\ast}}_{i^{\ast}}+R^{A}_{i^{\ast}}$. Then, the smallest amount of deposited token A in $\textbf{R}+a_{lp}^{\prime}$ is larger than that in $\textbf{R}+a_{lp}$:
@@
(2) The revenues of actions $a_{lp}$ and $a_{lp}^{\prime}$: To compare the revenue due to both actions, we consider the strategies and utility of the subsequent trader. Any best response of the trader $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ uses the smallest pool (Lemma G-A3):
@@
Thus, the revenue following action $a_{lp}^{\prime}$ is (using Equation VI-A4):
@@
Interpreting Equation G-B3, the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ is larger than the smallest pool in $\textbf{R}+a_{lp}$:
@@
From lemma G-B1, the trading fee of trading in a larger pool is larger,
$\textit{tf}(R^{A}_{i^{\ast}}+l^{A^{\prime}}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+l^{A^{\prime}}_{i^{\ast}}),b^{\textit{BA}})>\\ \textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\,\,.$ (113)
From the definition of $i^{\ast}$ in Equation G-B3, we have
@@
Combine Equations 112 and 114, we have
@@
Combining Equations 109110113 and 115, the revenue of the liquidity provider with action $a_{lp}^{\prime}$ is higher than with action $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$:
@@
Combining Equations 107 and 108, the utility of the liquidity provider under action $a_{lp}^{\prime}$ is higher than that under action $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$:
$U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})>\\ U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (117)
@@
$U_{lp}(\textbf{R},\pi_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})=\\ \sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\left(\begin{subarray}{c}\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\\ \times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\end{subarray}\right)\,\,.$ (119)
@@
$=$ $\sum_{a_{lp}^{\prime}\in\mathcal{A}_{lp}(l^{A},l^{B})}\left(\begin{subarray}{c}\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},a_{lp}^{\prime})\\ \times U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})\end{subarray}\right)$
@@
Combining Equations 119  G-B3121 and 122, the utility of the liquidity provider under $\pi_{lp}^{\prime}$ is higher than that under $\pi_{lp}$:
$U_{lp}(\textbf{R},\pi_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})>\\ U_{lp}(\textbf{R},\pi_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$